# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v2

@@ IV-C1 MSE and Huber Loss
Possible choices for the loss function $\mathcal{L}(\cdot)$ include MSE and Huber loss [37]. We explain these as follows.
@@
Figure 4: 5G NR time-frequency resource block with Type A DM-RS.
@@ IV-C2 Complexity
The proposed A-MMSE is fundamentally distinguished from existing DNN-based channel estimation approaches. To be specific, the proposed A-MMSE method learns a linear filter matrix $\mathbf{W}_{\text{A-MMSE}}$ during a dedicated training phase. As a result, once trained, channel estimation reduces to a single matrix-vector multiplication, as shown in (19). This enables highly efficient inference, with minimal computational cost and requiring only $2(NM\times L)$ real-valued parameters to store $\mathbf{W}_{\text{A-MMSE}}$. In contrast, as discussed earlier, conventional DNN-based methods [15, 9] learn a neural network that maps pilot inputs directly to channel outputs. As a result, each inference (i.e., channel estimation) involves propagating the input through multiple layers of nonlinear operations, leading to higher inference complexity. Specifically, the A-MMSE performs channel estimation using only linear operations with approximately 70,000 $\left(=NM\times L\right)$ parameters, whereas Channelformer requires about 110,000 parameters and includes nonlinear components, such as GELU activations and convolutional layers.
@@ VI-A Simulation Setup
For numerical simulations, we follow the 5G NR DM-RS configuration from 3GPP TS 38.211 [38]. DM-RS’s time-frequency density adapts to channel conditions: front-loaded configurations suit slowly varying channels, while rapidly varying channels require additional symbols to track the dynamics. Frequency selectivity further affects DM-RS pattern design, requiring a trade-off between estimation accuracy and pilot overhead. We illustrate this in Fig. 6.
@@
Due to space limitations, we focus on three representative TDL models: TDL-B (typical urban), TDL-D (high mobility), and TDL-E (mmWave), which together span a broad range of deployment scenarios across frequency bands. The diverse subcarrier spacings specified in Table I, corresponding to different symbol durations, are chosen intentionally to evaluate the proposed method across the fundamental trade-off between Doppler resolution and robustness to channel time-variations. As benchmark methods, we consider the following. For SP-based approaches, we include LS, MMSE, and 1D-MMSE. For DNN-based approaches, we include ChannelNet [9] and Channelformer [15]. Furthermore, in the final subsection, we also compare the performance of a sensing-aided approach [7] and A-MMSE.
@@ VII Conclusion
In this paper, we proposed the A-MMSE method for OFDM channel estimation. Our key idea is to integrate the classical SP-based framework with the learning capability of the DNN-based approach. Specifically, the proposed A-MMSE learns the linear filtering matrix via the Attention Transformer and applies it for channel estimation, achieving both high estimation accuracy and substantial computational efficiency. In particular, to enhance the learning of the linear filtering matrix, we developed a novel two-stage Attention encoder that effectively captures the frequency and temporal correlation structures of the OFDM channel. Building on this architecture, we further introduced a rank-adaptive extension of A-MMSE, referred to as the RA-A-MMSE. The RA-A-MMSE stands out by dynamically adjusting the rank of the linear filtering matrix, thereby controlling the computational overhead associated with channel estimation. Through extensive numerical simulations using 3GPP TDL channel models, we demonstrated that the proposed A-MMSE provides significant performance gains in channel estimation compared to conventional SP-based methods (e.g., LS and MMSE) as well as DNN-based methods (e.g., ChannelNet [9] and Channelformer [15]). We further showed that the RA-A-MMSE effectively balances estimation performance and computational complexity, offering substantial reductions in computational overhead with minimal performance degradation.
@@ Appendix A Proof of Kronecker Product Structure
Consider the time-varying channel impulse response $h(t,\tau)$ [39] with time-frequency transfer function
@@
Under WSSUS [39]: $\mathbb{E}\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]=r_{h}(t-t^{\prime};\tau)\delta(\tau-\tau^{\prime}).$ This yields 2D stationarity:
@@
Assuming separable scattering [35] $C_{H}(\tau,\nu)=P_{\tau}(\tau)\cdot S_{\nu}(\nu)$, we can substitute as

### v3

@@ IV-C1 MSE and Huber Loss
Possible choices for the loss function $\mathcal{L}(\cdot)$ include MSE and Huber loss [huber1964robust]. We explain these as follows.
@@
Figure 4: TDL-D: Channel estimation NMSE comparing Huber loss vs. MSE loss
@@ IV-C2 Complexity
The proposed A-MMSE is fundamentally distinguished from existing DNN-based channel estimation approaches. To be specific, the proposed A-MMSE method learns a linear filter matrix $\mathbf{W}_{\text{A-MMSE}}$ during a dedicated training phase. As a result, once trained, channel estimation reduces to a single matrix-vector multiplication, as shown in (19). This enables highly efficient inference, with minimal computational cost and requiring only $2(NM\times L)$ real-valued parameters to store $\mathbf{W}_{\text{A-MMSE}}$. In contrast, as discussed earlier, conventional DNN-based methods [luan2023channelformer, soltani:commlett:19] learn a neural network that maps pilot inputs directly to channel outputs. As a result, each inference (i.e., channel estimation) involves propagating the input through multiple layers of nonlinear operations, leading to higher inference complexity. Specifically, the A-MMSE performs channel estimation using only linear operations with approximately 70,000 $\left(=NM\times L\right)$ parameters, whereas Channelformer requires about 110,000 parameters and includes nonlinear components, such as GELU activations and convolutional layers.
@@ VI-A Simulation Setup
For numerical simulations, we follow the 5G NR DM-RS configuration from 3GPP TS 38.211 [3GPP-TS-38.211]. DM-RS’s time-frequency density adapts to channel conditions: front-loaded configurations suit slowly varying channels, while rapidly varying channels require additional symbols to track the dynamics. Frequency selectivity further affects DM-RS pattern design, requiring a trade-off between estimation accuracy and pilot overhead. We illustrate this in Fig. 6.
@@
Due to space limitations, we focus on three representative TDL models: TDL-B (typical urban), TDL-D (high mobility), and TDL-E (mmWave), which together span a broad range of deployment scenarios across frequency bands. The diverse subcarrier spacings specified in Table I, corresponding to different symbol durations, are chosen intentionally to evaluate the proposed method across the fundamental trade-off between Doppler resolution and robustness to channel time-variations. As benchmark methods, we consider the following. For SP-based approaches, we include LS, MMSE, and 1D-MMSE. For DNN-based approaches, we include ChannelNet [soltani:commlett:19] and Channelformer [luan2023channelformer]. Furthermore, in the final subsection, we also compare the performance of a sensing-aided approach [Qing2024IoTJ] and A-MMSE.
@@ VII Conclusion
In this paper, we proposed the A-MMSE method for OFDM channel estimation. Our key idea is to integrate the classical SP-based framework with the learning capability of the DNN-based approach. Specifically, the proposed A-MMSE learns the linear filtering matrix via the Attention Transformer and applies it for channel estimation, achieving both high estimation accuracy and substantial computational efficiency. In particular, to enhance the learning of the linear filtering matrix, we developed a novel two-stage Attention encoder that effectively captures the frequency and temporal correlation structures of the OFDM channel. Building on this architecture, we further introduced a rank-adaptive extension of A-MMSE, referred to as the RA-A-MMSE. The RA-A-MMSE stands out by dynamically adjusting the rank of the linear filtering matrix, thereby controlling the computational overhead associated with channel estimation. Through extensive numerical simulations using 3GPP TDL channel models, we demonstrated that the proposed A-MMSE provides significant performance gains in channel estimation compared to conventional SP-based methods (e.g., LS and MMSE) as well as DNN-based methods (e.g., ChannelNet [soltani:commlett:19] and Channelformer [luan2023channelformer]). We further showed that the RA-A-MMSE effectively balances estimation performance and computational complexity, offering substantial reductions in computational overhead with minimal performance degradation.
@@ Appendix A Proof of Kronecker Product Structure
Consider the time-varying channel impulse response $h(t,\tau)$ [Bello1963] with time-frequency transfer function
@@
Under WSSUS [Bello1963]: $\mathbb{E}\left[h(t,\tau)h^{*}(t^{\prime},\tau^{\prime})\right]=r_{h}(t-t^{\prime};\tau)\delta(\tau-\tau^{\prime}).$ This yields 2D stationarity:
@@
Assuming separable scattering [Yoojin20152DLMMSE] $C_{H}(\tau,\nu)=P_{\tau}(\tau)\cdot S_{\nu}(\nu)$, we can substitute as